OpenStudy (anonymous):

Alex ran 2 7/10 ( mixed number ) miles on Monday. He ran twice as far on Tuesday. How far did Alex run in all those two days?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2\frac{ 7 }{ 10 }+2\left( 2\frac{ 7 }{ 10 } \right)=?\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\[2\frac{ 7 }{ 10 }=\frac{ 27 }{ 10 }\]

OpenStudy (anonymous):

you're welcome